OpenStudy (anonymous):

a. Sarah and her uncle, Takota went fishing in the Grand River, and each caught one fish. Sarah's fish was 5/8 of a foot long and Takota's was 2/3 of out a foot long. Which Fish is longer? Explain

OpenStudy (anonymous):

2/3 is bigger. Because the portions are bigger and there are 2 out of the 3 that would be needed.

OpenStudy (anonymous):

5/8 = 15 / 24 2/3 = 16/ 24